OpenStudy (anonymous):

what would 2pi/pi/4 simplified be for a period

OpenStudy (anonymous):

\[\frac{2\pi}{\frac{\pi}{4}}=2\pi\times \frac{4}{\pi}\] better known as \(8\)

OpenStudy (anonymous):

so pi/4 =8 ?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\(\frac{\pi}{4}\neq 8\)

OpenStudy (anonymous):

two pi divided by pi over 4 is 8 just like two divided by one fourth is 8

OpenStudy (anonymous):

i just need pi/4 simplified

OpenStudy (anonymous):

the 2pi part is for the period